40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 321 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  714   Tue Jul 22 13:15:14 2008 robUpdatePSLNote from R. Abbott re: the PMC

Quote:
an email from Rich:
Your PZT is broken.

R


Quelle surprise

Frown
  713   Tue Jul 22 11:55:22 2008 ranaUpdatePSLNote from R. Abbott re: the PMC
an email from Rich:
Your PZT is broken.

R
  712   Tue Jul 22 09:24:17 2008 steveUpdatePSLlaser power
Laser power reality of 120 days
Attachment 1: power120d.jpg
power120d.jpg
  711   Tue Jul 22 03:03:22 2008 John, RobUpdatePSLFSS open loop transfer function
With the common gain slider maxed out the unity gain frequency is 58kHz.

The reference cavity refl diode appears to be okay. RF OUT/ TEST IN transfer function was normal.
There is a ~220mV offset in the RF out. We removed this using a coupler - no change. We also checked the
diode->FSS cable.

Tomorrow I'll take a closer look at the board.
  710   Mon Jul 21 19:55:16 2008 rana,jenneConfigurationIOONoise in MC_F
Jenne put the MC board on extender today - its still that way but everything is probably connected (check AO).

We measured the TFs of the DAQ section for MC_F because of how everything looked wrong in the plots Jenne
put in the log earlier. Everything we measured today seemed to jive with the schematic. We also looked up
the original traveler for this board which Betsy filled in years ago: it also is in spec for the DAQ filtering.

So then I looked at the power spectrum of the output signal to the VCO. It had lots of HFC (high frequency crap).
I adjusted the parameters of the FSS (common gain, fast gain, RF phase Astonished) and lowered the MC common gain. This
produced a global minimum in that 4D parameter space.

I think that basically, the FSS gain is too low even with the common gain slider maxed. Having the fast gain up
at 19 dB like I had left it was bad - even though it minimizes the PC control signal, it produces a lot of HFC
up around 100 kHz in MC_F. After John (finally) gets around to measuring the FSS loop we can figure out how to
better tune this. The MC gain then has to be tuned so as to best minimize the HFC given the new FSS gain; there's
basically no coupling from the MC gain to the FSS loop shape so its always best to tune the FSS first. Pleased

The RF phase of the FSS was a mystery - I have no idea why it should do anything and I have never heard of this
and I don't know why I tried it today. But...changing it by ~0.6-0.7 slider units reduced the HFC by another factor
of ~3. Somebody should put this slider into units of degrees.8-)

Here's a table of the changes. Please make these the new nominals:
you asked for:   diff 2008/07/21,13:00 2008/07/22,2:44:16 utc '.*FSS.*|.*MC.*'
LIGO controls: differences, 2008 07/21 13:00:00 utc vs. 2008 07/22 02:44:16 utc
__Epics_Channel_Name______   __Description__________   __value1____     __value2____
C1:IOO-MC_REFL_GAIN                                    22.000000        19.000000
C1:IOO-MC_REFL_OFFSET                                  0.818380         0.818100
C1:PSL-FSS_MGAIN                                       10.000000        30.000000
C1:PSL-FSS_PHCON                                       2.073170         1.413170

The attached plot shows the "SERVO" TNC output of the board; this is supposed to be the same as the voltage going to the
VCO box. So its V/Hz transfer function is flat above 40 Hz. Tomorrow Jenne will post more data and remove the extender
board.

Since I only used an SR785, I only saw noise up to 100 kHz. Its key to use an RF spectrum analyzer when checking out
the FSS and the MC systems.
Attachment 1: SCRN0024.GIF
SCRN0024.GIF
  709   Mon Jul 21 19:48:57 2008 SharonUpdate how tp restart C1ASS
How to restart C1ASS:

1. reboot
2. as root: caltech/target/c1ass:> ./startass
3. no need for root: burtgooey

that's it...
  708   Mon Jul 21 15:52:22 2008 steveBureaucracySAFETYfire alarms test
The fire alarm test and evacuation drill was successfully completed at 13:45 Wednesday, July 16, 2008

Everybody was on time for the 40m meeting.
  707   Mon Jul 21 14:26:11 2008 MaxSummaryPEMAdded Channels
The following channels have been added.

Channel Name DAQ port
C1 : PEM-MAG_BSC_X 27
C1 : PEM-MAG_BSC_Y 28
C1 : PEM-MAG_BSC_Z 29

Jenne and I ran the wires from near the beam splitter chamber (as described in a previous elog) to the rack Y7 and plugged the labeled BNC's into ports 27-29. The computer was c0dcu1. John then restarted the frame builder and Alberto and I restarted the front end of c0dcu1 as per the wiki's instructions. The channels seem to be working. - Max.
  706   Mon Jul 21 11:54:00 2008 JenneUpdateGeneralMC Servo Board
I pulled the MC Servo Board again, to check the components that are on the board, and compare them with the schematics. The filters that I'm interested in on the Fast Path haven't been changed. The high pass filters on the Fast Path have been changed.
Component      Schematic      Actual
---------      ---------      ------
C140           10u            open
C144           10u            open
C149           open           a gray Cap.  value unknown
C141           10u            open
C145           10u            open
R97            1.58K          0
R99            open           1130
R103           open           1130
R100           open           0
R104           100            1130
R98            1.58K          open
R109           367            365

Board is back in, and MC locks.
  705   Mon Jul 21 10:00:42 2008 steveUpdatePEMifo room temp
Five years of temp trend with 5 C fluctuation. The campus chilled water supply was upgraded ~ 1 year ago.

I requested our thermostats to be calibrated.
Attachment 1: temp5y.jpg
temp5y.jpg
  704   Mon Jul 21 09:52:05 2008 SharonUpdate Adaptive code changes
Thanks to Alex, we now save the coefficients of the adaptive filter every cycle. When we choose ENABLE: OFF on the MEDM screen, suppressing the signal to the MC1, we stop and save the last coefficients. When enabling it again, it starts from the last coefficients saved. I will take some measurements today to check how this contributes to the adaptation rate. If you change the number of taps or the number of AUX channels, the coefficients are again set to zero.
  703   Sat Jul 19 19:41:56 2008 YoichiAoGPSLThe author of the entry 702 is Yoichi not Rob
I made a mistake.
  702   Sat Jul 19 19:39:44 2008 robUpdatePSLPMC PZT investigation

Quote:

Quote:
The 64k is the series resistor after the PA85, not the modeled resistance of the PZT itself.

Yes. What I meant was that because the measured voltage across the PZT was a half of the open voltage of the HV amp, the DC impedance of the PZT is expected to be similar to the output impedance of the HV amp. Of course, I don't think the DC impedance of a normal PZT should be such low.
I'm puzzled by the discrepancy between this expected DC impedance and the directly measured impedance by the Fluke DVM (1.5M Ohm).
One possibility is that the PZT leaks current only when a high voltage is applied.
  701   Fri Jul 18 23:24:24 2008 robUpdatePSLPMC PZT investigation

Quote:
I measured the HV coming to the PMC PZT by plugging it off from the PZT and hooking it up to a DVM.
The reading of DVM is pretty much consistent with the reading on EPICS. I got 287V on the DVM when the EPICS says 290V.

Then I used a T to monitor the same voltage while it is connected to the PZT. I attached a plot of the actual voltage measured by the DVM vs the EPICS reading.
It shows a hysteresis.
Also the actual voltage drops by more than a half when the PZT is connected. The output impedance of the HV amp is 64k (according to the schematic). If I believe this number, the impedance of the PZT should also be 64k. The current flowing the PZT is 1.6mA at 200V EPICS reading.
The impedance of the PZT directly measured by the DVM is 1.5M ohm, which is significantly different from the value expected above. I will check the actual output impedance of the HV amp later.
The capacitance of the PZT measured by the DVM is 300nF. I don't know if I can believe the DVM's ability to measure C.

I noticed that when a high voltage is applied, the actual voltage across the PZT shows a decay.
The second plot shows the step response of the actual voltage.
The voltage coming to the PZT was T-ed and reduced by a factor of 30 using a high impedance voltage divider to be recorded by an ADC.
The PMCTRANSPD channel is temporarily used to monitor this signal.
After the voltage applied to the PZT was increased abruptly (to ~230V), the actual voltage starts to exponentially decrease.
When the HV was reduced to ~30V, the actual voltage goes up. This behavior explains the weird exponential motion of the PZT feedback signal when the PMC is locked.
The cause of the actual voltage drop is not understood yet.
From the above measurements, we can almost certainly conclude that the problem of the PMC is in the PZT, not in the HV amp nor the read back.


I'd believe the Fluke's measurement of capacitance. Here's some info from PK about the PZT:


Quote:

But the PMC ones were something like
0.750 in. thick x 0.287 in. thick. 2 microns per 200 V displacement,
resonant frequency greater than 65 kHz. Typical capacitance is around 0.66
uF.


If the PZT capacitance has dropped by a factor of two, that seems like a bad sign. I don't know what to expect for a resistance value of the PZT, but I wouldn't be surprised if it's non-Ohmic. The 64k is the series resistor after the PA85, not the modeled resistance of the PZT itself.
  700   Fri Jul 18 19:43:55 2008 YoichiDAQComputersPSL fast channels cannot be read by dataviewer
At this moment only the PSL fast channels have trouble.
Rob restarted fb40m, c1IOVME, but no effect.
  699   Fri Jul 18 19:41:09 2008 YoichiUpdatePSLPMC PZT investigation
I measured the HV coming to the PMC PZT by plugging it off from the PZT and hooking it up to a DVM.
The reading of DVM is pretty much consistent with the reading on EPICS. I got 287V on the DVM when the EPICS says 290V.

Then I used a T to monitor the same voltage while it is connected to the PZT. I attached a plot of the actual voltage measured by the DVM vs the EPICS reading.
It shows a hysteresis.
Also the actual voltage drops by more than a half when the PZT is connected. The output impedance of the HV amp is 64k (according to the schematic). If I believe this number, the impedance of the PZT should also be 64k. The current flowing the PZT is 1.6mA at 200V EPICS reading.
The impedance of the PZT directly measured by the DVM is 1.5M ohm, which is significantly different from the value expected above. I will check the actual output impedance of the HV amp later.
The capacitance of the PZT measured by the DVM is 300nF. I don't know if I can believe the DVM's ability to measure C.

I noticed that when a high voltage is applied, the actual voltage across the PZT shows a decay.
The second plot shows the step response of the actual voltage.
The voltage coming to the PZT was T-ed and reduced by a factor of 30 using a high impedance voltage divider to be recorded by an ADC.
The PMCTRANSPD channel is temporarily used to monitor this signal.
After the voltage applied to the PZT was increased abruptly (to ~230V), the actual voltage starts to exponentially decrease.
When the HV was reduced to ~30V, the actual voltage goes up. This behavior explains the weird exponential motion of the PZT feedback signal when the PMC is locked.
The cause of the actual voltage drop is not understood yet.
From the above measurements, we can almost certainly conclude that the problem of the PMC is in the PZT, not in the HV amp nor the read back.
Attachment 1: Hysteresis.png
Hysteresis.png
Attachment 2: StepResponse.png
StepResponse.png
  698   Fri Jul 18 19:30:20 2008 MashaUpdateAuxiliary lockingmoving from 40m
I will be working in the basement of Bridge probably starting next week; I moved the NPRO laser and some of the optics from my mach zehnder setup on the SP table to Bridge. Thanks for your help!
  697   Fri Jul 18 19:15:15 2008 JenneUpdateIOOChecking out the MC Servo Board

Quote:
[Jenne, Max]

I haven't checked to ensure that the MC still locks, because Yoichi is doing stuff on the PSL table, but I didn't change anything on the board, and hooked all the cables back where they were, so hopefully it's all okay.


I put the PMC back and the MC now locks.
  696   Fri Jul 18 17:12:35 2008 JenneUpdateIOOChecking out the MC Servo Board
[Jenne, Max]

One of the things that Rana thinks that might be causing my MC_F calibration to be off is that the MC Servo Board's filters don't match those on the schematics. Max and I pulled the MC servo board today to check resistor and capacitor values. Alberto needed the Mode Cleaner, so we put the board back before finishing checking values. We will probably pull the board again next week to finish checking the values.

I haven't checked to ensure that the MC still locks, because Yoichi is doing stuff on the PSL table, but I didn't change anything on the board, and hooked all the cables back where they were, so hopefully it's all okay.
  695   Fri Jul 18 17:06:20 2008 JenneUpdateComputersComputers down for most of the day, but back up now
[Sharon, Alex, Rob, Alberto, Jenne]

Sharon and I have been having trouble with the C1ASS computer the past couple of days. She has been corresponding with Alex, who has been rebooting the computers for us. At some point this afternoon, as a result of this work, or other stuff (I'm not totally sure which) about half of the computers' status lights on the MEDM screen were red. Alberto and Sharon spoke to Alex, who then fixed all of them except C1ASC. Alberto and I couldn't telnet into C1ASC to follow the restart procedures on the Wiki, so Rob helped us hook up a monitor and keyboard to the computer and restart it the old fashioned way.

It seems like C1ASC has some confusion as to what its IP address is, or some other computer is now using C1ASC's IP address.

As of now, all the computers are back up.
  694   Fri Jul 18 16:57:37 2008 JenneUpdateIOOCalibrated MC_F
I have calibrated MC_F. The conversion factor is 137.49 MHz/count.

The calibration data taken is attached, along with a calibrated power spectrum.

On the data plot, the x axis is volts from the C1:IOO-MC_FAST_MON channel, with the calibration between FAST_MON and MC_F = -788.18 volts/count.
The linear term of the fit line = -0.085MHz/volt. Error bars are +/- 1 in the last digit of what the spectrum analyzer gave me for frequency (+/- 0.01MHz).

The net conversion factor is then (-788.18)*(-0.085)*(2) = 137.49 MHz/count. The factor of 2 is because the light passes through the AOM twice.

On the power spectrum,
REF0 and REF1 = MC unlocked, HEPAs on, MC Refl gain = 22
REF2 and REF3 = MC locked, HEPAs on, MC Refl gain = 22
REF4 and REF5 = MC locked, HEPAs on, MC Refl gain = 19
REF6 and REF7 = MC locked, HEPAs off, MC Refl gain = 19
Attachment 1: MC_Fcalib.png
MC_Fcalib.png
Attachment 2: 20080717MC_F-MC_I.pdf
20080717MC_F-MC_I.pdf
  693   Fri Jul 18 12:24:15 2008 josephb, EricConfigurationCamerasChanged Lenses on GC750 at ETMX
We removed the giant TV zoom lens and replaced it with a much smaller fixed zoom lens. Currently it views the entire optic. We have another (also small) zoom lens which focuses much better on the spot itself. With how far back the camera is currently placed, neither of these fixed zoom lenses can touch or hit the view port or the chamber while still attached to camera and mount, even using all of the mount's motion range. So this should be less of a safety issue.

Ideally, we'd like to get some images of the full optic (including osems and so forth) with the X-arm locked, and then use the higher zoom lens while still locked, to get images we can use to calibrate the x and y length scales.
  692   Thu Jul 17 20:13:34 2008 YoichiUpdatePSLPMC alignment/mode matching effort
I'm working to improve the mode matching of PMC.
Because I noticed that the beam was hitting the aperture of the EOM for PMC, I moved the EOM a little bit to maximize the transmission.
This did not change the alignment to the reference cavity but changed the alignment of the PMC a lot.
So I adjusted it back.
The alignment of the PMC can be easily optimized but the Hermite 02 mode still remains. This means the mode matching is bad.
Moving the lenses by a small amount (a few mm) did not change the height of 02 mode.
I'm planning to move the lenses by a large amount tomorrow. But it will destroy the alignment to the PMC.
So I installed two irises in the beam path after the lenses to remember the alignment roughly.
Right now the PMC transmission is slightly worse than before because the lens positions are not good.
  691   Thu Jul 17 16:39:58 2008 Max JonesUpdateDAQMagnetometer Installed
Today I installed the magnetometer near the beam splitter chamber. It is located on the BSC chamber at head height on the inner part of the interferometer (meaning I had to crawl under the arms to install it). I don't think I disturbed anything during installation but I think that's it's probably prudent to tell everyone that I was back there just in case. I plan to run 3 BNC cables (one for each axis) from the magnetometer to the DAQ input either tonight or tomorrow. Suggestions are appreciated. - Max.
  690   Thu Jul 17 13:08:37 2008 JohnSummaryLSCHOM resonances in the arms
On Tuesday night we attempted to lock the full DRFPMI in the optical spring configuration with the +f2 sideband resonant in the SRC.
Despite having no problems locking on the +f2 in a DRM we couldn't lock the full IFO.

There was some discussion about whether a +f2 higher order mode resonance in the arms could cause this problem.

I calculated the positions of the first six higher order modes for the carrier and all sidebands (using Siegman p 762 (23) with a plus sign).
Plot attached. Colors indicate different frequency components, numbers are the mode index (m+n). Thick lines are fundamental modes of
the sidebands. Heights of HOM indicators have been scaled by 1/(m+n)^2.

It appears that the first order transverse mode of the +166 is indeed partially resonant. We might try to tweak the sideband frequencies a
little to see if this helps us. It would probably be prudent to measure the MC length first.

Numbers used:

L = 38.5750; %average of Alberto's recent measurements elog #556
Retm = 57.375;
f166 = 165.977195e6;
f33 = 33.195439e6;
Attachment 1: HOMresonances.png
HOMresonances.png
  689   Thu Jul 17 12:15:21 2008 EricUpdatePSLSwept PMC PZT voltage range
I unlocked the PMC and swept over C1:PSL-PMC_RAMP's full range a couple of times this morning.  The PMC should now be relocked and returned 
to normal.
  688   Thu Jul 17 08:30:15 2008 steveUpdatePSLPMC relocked manually
The PMC pzt HV and the servo gain adj. are railing at max this morning

Why is it on the decreasing side of FSS_RTTEMP slope?
Attachment 1: pmc4d.jpg
pmc4d.jpg
  687   Thu Jul 17 00:59:18 2008 JenneSummaryGeneralFunny signal coming out of VCO
While working on calibrating the MC_F signal, Rana and I noticed a funny signal coming out of the VCO. We expect the output to be a nice sine wave at about 80MHz. What we see is the 80MHz signal plus higher harmonics. The reason behind the craziness is to be determined. For now, here's what the signal looks like, in both time and frequency domains.

The first plot is a regular screen capture of a 'scope. The second is the output of the SR spectrum analyzer, as seen on a 'scope screen. The leftmost tall peak is the 80MHz peak, and the others are the harmonics.
Attachment 1: VCOout_time.PNG
VCOout_time.PNG
Attachment 2: VCOout_freq.PNG
VCOout_freq.PNG
  686   Wed Jul 16 22:29:05 2008 MashaUpdateAuxiliary lockinglong measurement

Quote:
I'm taking a measurement on the SR785 spectrum analyzer at low frequencies, so I'm going to leave it by the symmetric port table for a while. Please don't move it!


all done thanks.
  685   Wed Jul 16 17:51:58 2008 MashaUpdateAuxiliary lockinglong measurement
I'm taking a measurement on the SR785 spectrum analyzer at low frequencies, so I'm going to leave it by the symmetric port table for a while. Please don't move it!
  684   Wed Jul 16 17:36:51 2008 JohnDAQPSLFSS input offset
I changed the nominal FSS input offset to 0 from 0.3. Tolerance remains unchanged at +/-0.05.
  683   Wed Jul 16 16:59:07 2008 AlbertoUpdateGeneralAligment
I think the two beams are aligned again - they both pass the Faraday, they match at the irises and all along the optical path on the AP table. Although the NPRO beam does not show up at the AS port.
  682   Wed Jul 16 16:28:14 2008 josephbConfigurationComputersFixed IP address on Switch
Realized today that the change I made back on June 30th to the switch was to the wrong switch. I had disabled the DHCP setting and mislabeled the switch in the control room (which seems to not have affected anything).

I've turned DHCP back on and labeled it correctly using the Netgear "Smartwizard discovery" program.
  681   Wed Jul 16 15:59:04 2008 josephb, EricConfigurationCamerasPMC trans camera path
In order to reduce saturation, we placed a Y1 plate (spare from the SP table) in transmission just before the GC650 camera looking at the PMC transmision. The reflection (most of the light) was dumped to a convient razor blade dump. We also removed the 0.3 and 0.5 ND filters and placed them in the 24 hour loan ND filter box.

Good exposure values to view are now around 3000 for that camera.
  680   Wed Jul 16 11:26:47 2008 Max JonesUpdate This Week
Baffles.

I got a battery for the magnetometer today which is slightly too large (~2 mm) in one dimension. Not sure what I'm going to do.

I'm attempting to calibrate the magnetometer but I'm having a hard time calibrating the axis that I cannot simply put through a coil parallel to the coils length. I have attempted to use the end fields of the solenoid but the measurements from the magnetometer are significantly different from the theoretical calculations.

I would appreciate suggestions. - Max.
  679   Wed Jul 16 11:00:15 2008 MashaUpdateAuxiliary lockingimproving ADC input for the mach zehnder setup and completely unrelated happenings
For most of last week, the SURFs + Jenne were helping Mike and Ken with "stray light control for
Enhanced LIGO", i.e. cleaning and baking many many baffles which will catch scattered light in the
interferometer.

Otherwise, the two channels of the Mach Zehnder which will be used to measure fibre noise were
balanced, which should reduce the effect of laser amplitude noise in phase noise detection. I have
set up two digital channels to collect time series data from the two photodiodes and took some
preliminary noise measurements. I will be using Matlab to combine the signals as to directly measure
the phase noise, and I wrote some Matlab code to speed up this process: loading the files,
manipulating time series data, and converting into frequency domain. Currently I am building a
filter that will attenuate the signal at frequencies below 1Hz and amplify at higher frequencies in
order to whiten the spectra and reduce ADC noise.
  678   Wed Jul 16 10:50:55 2008 EricSummaryCamerasWeekly Summary
Finished unwrapping, cleaning, baking, wrapping, wrapping again, packing, and shipping the baffles.

Attempted to set up the Snap software so that it could talk directly to EPICS channels. This is not currently working due to a series of very strange bugs in compiling and linking the channel access libraries. Alex Ivanov directed Joe and me to a script and makefile that are similar to what we're trying to do and it may solve our problem, but at the moment this still doesn't work. We're currently using a workaround that involves making unix system calls to ezca command line tools, but this is too hacky to leave in the final program.

Attempted to fit Josh's PZT voltage vs power plot of the OMC (from about a year ago) to lorentzians in order to try to develop fitting tools for more recent data. This isn't working, due to systematic error distorting the shapes of the peaks. Good fits can be obtained by cutting the number of points to a very small number around the peak of resonance, but this leads to such a small percentage of the peak being used that I don't trust these results either. (In the graph (shows the very top of the tallest peak): blue is Josh's original data, green is a fit to this peak using the top 66% of the peak and arbitrary, equal values for the error on each point, red is Josh's data averaged over bins of size 0.005, teal is a fit to these bins where the error on each point is the standard deviation of each bin, and magenta is a fit to these same bins, except cropped to the top ~10% of the peak, x-axis is voltage, y-axis is transmission power). Rana suggested that I take my own sweeps of the PMC using scripts that are already written: I'm currently figuring out where these scripts are and how to use them without accidently breaking something.

We've begun running the Snap software for long periods of time to see how stable it is. Currently, its only problem appears to be that it memory leaks somewhat: it was up 78% memory usage after a little over an hour. It doesn't put much strain on the computer, using only ~20% CPU. Stress put on the network from the constant transfer of images from the camera to the computer is not yet known.
Attachment 1: AttemptedPeakFit3.tiff
  677   Wed Jul 16 09:27:17 2008 steveUpdateALARMPSL-FSS_RMTEMP alarm is false
Morning alarm sound is good for people who does not drink coffee.
Our 40m alarm is on every morning.
Those whom are not here in the morning thinks that this beeping sound is inspirational.
Would someone change this sound into less punishing form, like mockingbird chirp....

The C1PSL_SETTINGS.adl (40mm PSL Settings ) indicating that
C1:PSL-FSS_INOFFSET (Input Offset Adjust ) should be 0.3 +-0.05 V (red warning tag )

Alarm Handler: 40M pointing to yellow grade warning of PSL-FSS_RMTEM
This is a false alarm.

Two years trend of these channels are here:
Attachment 1: frmtemp2y.jpg
frmtemp2y.jpg
  676   Tue Jul 15 19:15:57 2008 ranaSummarySUSETMX Dewhitening characterization
Since the boys found that the ETM dewhitening transient was kicking the IFO out of lock we
decided to investigate.

First, we wrote a script to diagnose and then tune the DC gain of the dewhitening filters'
digital compensation filter (a.k.a. FM9 or SimDW). It is in the scripts/SUS/ directory
and is called dwgaintuner. It puts in an offset on each coil's DAC channel and
then reads back the Vmon on the coil driver with the DWF on and off. It reports the ratio of these
voltages which you can then type into the FM9/SimDW filter's gain field. We learned that the
difference between the analog DWF path and the bypass path was ~3% (which is consistent with
what you expect from the use of 1% resistors). We need to repeat this for all of the rest of
the suspended optics except for MC1 and MC3.

This Vmon method is better than what's used at the sites so we will export this new technology.

The attached plot shows some switching transients with only the local damping on:
BLUE:   Output of filter bank during an FM9 turn off. This is the transient which goes to the DAC.
        The transients are mostly of the same magnitude as this.
RED:    This is the input of the filter module during another such transient.
GREEN:  Tried another switch; this time I filtered the time series in DTT by typing the SimDW into
        the Triggered Time Series filter field. This should be simulating what comes out of the
        output of the DW board - to convert to volts multiply by 15/32768.
PURPLE: Same kind of filtering as the GREEN, but with also a double 30 Hz highpass to remove the low
        frequency damping control signals. You can see that the total transient is only ~5 counts
        or ~1 mV at the coil driver output. This is comparable to the relative offset in the bypass
        and filter paths.
Attachment 1: etmx-dw-transient.pdf
etmx-dw-transient.pdf
  675   Tue Jul 15 12:44:08 2008 JohnSummaryLockingDRFPMI with DC readout
Rob, John

Last night, despite suspect alignment, we were again able to reduce the CARM offset to zero using
the RF signal.We were also able to transfer to dc readout taking calibrated spectra in both states.
DC readout shows a marked improvement over RF above ~1kHz but introduces some noise around 100Hz.
Broadband sensitivity appears to be more than ten times worse than previously. The calibration
being used remains to be confirmed.

Engaging the ETMY dewhitening caused lock to be lost. We'll check this today. The OMC alignment loops
may also need some attention.

We looked at REFL_166 as a candidate for CARM, at present POX still looks better.

The DARM filters were modified to reduce excess noise around 3Hz. Updating filter coefficients does
not cause loss of lock.
  674   Tue Jul 15 12:23:22 2008 JenneUpdateGeneralMC2 Watchdog tripped
Alberto, Jenne

Mode Cleaner was unlocked. We checked, and found that MC2's watchdog was tripped. It didn't look like anything bad was going on, so we turned the optic back on, and tried to relock the MC. It looks like the Mode Cleaner is now locked, but the lock bit on the LockMC screen is still red. I don't know what's up.
  673   Tue Jul 15 11:47:56 2008 JenneDAQPEMAccelerometer channels in ASS Adapt MEDM screen
Jenne, Sharon

We have traced which accelerometers correspond to which channels in the C1ASS_TOP MEDM screen.

Accelerometer Channel
------------- --------------------------
MC1-X C1:ASS-TOP_PEM_2_ADAPT_IN1
MC1-Y C1:ASS-TOP_PEM_3_ADAPT_IN1
MC1-Z C1:ASS-TOP_PEM_4_ADAPT_IN1
MC2-X C1:ASS-TOP_PEM_5_ADAPT_IN1
MC2-Y C1:ASS-TOP_PEM_6_ADAPT_IN1
MC2-Z C1:ASS-TOP_PEM_7_ADAPT_IN1

SEISMOMETER C1:ASS-TOP_PEM_1_ADAPT_IN1
  672   Tue Jul 15 10:24:57 2008 steveUpdatePSLPMC temp & pzt voltage
The PMC pzt HV was happy with no HEPA temp stability.
Can we thermally insulate the pmc ?
Attachment 1: pmctemp.jpg
pmctemp.jpg
  671   Tue Jul 15 10:09:42 2008 EricDAQCamerasDid anyone kill the picture taking process on Mafalda?
Did anyone kill the process on Mafalda that was taking pictures of the end mirror of the x-arm last Friday? I need to know whether or not it crashed of its own accord.
  670   Tue Jul 15 09:47:09 2008 steveUpdatePEMlab temps and particles
All air condtion units were serviced last Friday.
AC filters are trying to control our particle counts but they have no capacity to match bad Pasadena conditions.
IFO room filters at CES were really clean.
Air make up filters inside and outside were dirty.
They showed the construction effect.
Control room and clean assembly units needed all filters replaced.

Note: the PSL-FSS_RCTEMP droped o.1C when enclouser HEPAs were turned back on
The RC temp controller should be better than that!
Attachment 1: temps24d.jpg
temps24d.jpg
  669   Mon Jul 14 21:34:10 2008 AlbertoUpdateGeneralabs cavity length measurement experiment

Quote:
Lately I've been dealing with the alignment of the interferometer to have a good beam spot at the AS port. Today the alignment script kept failing because of computer problems (failure of the frame builder) and also because the IFO was probably too far from the range where the automatic alignment works.

An other problem I keep having with the alignment of the optics on the AP table is with multiple reflection beams of the NPRO beam at the Farady.
Although I believe that now the two beams are quite well aligned, I don't see any reflection of the secondary beam from the IFO anymore.

It's like the more I try to improve the alignment, the worse I get from the beam matching. I'll keep working on this.


Realigning the OSA I also had to move a little bit the mirror that reflects the IFO beam of at the AS port in order to raise the beam height. This had the effect of changing the position of the AS spot on the camera and on the monitors.

Tonight with John, we made sure that the AS beam was still aligned to the PD.
  668   Mon Jul 14 19:15:43 2008 AlbertoUpdateGeneralabs cavity length measurement experiment
Lately I've been dealing with the alignment of the interferometer to have a good beam spot at the AS port. Today the alignment script kept failing because of computer problems (failure of the frame builder) and also because the IFO was probably too far from the range where the automatic alignment works.

An other problem I keep having with the alignment of the optics on the AP table is with multiple reflection beams of the NPRO beam at the Farady.
Although I believe that now the two beams are quite well aligned, I don't see any reflection of the secondary beam from the IFO anymore.

It's like the more I try to improve the alignment, the worse I get from the beam matching. I'll keep working on this.
  667   Mon Jul 14 12:43:07 2008 JohnSummaryComputersRestarted fb40m, tpman and c1ass
  666   Mon Jul 14 10:57:00 2008 KojiFrogsEnvironmentSomeone at 40M sent LHO water of life
Someone at the 40m sent Mike@LHO a pound of peets coffee with the name of Koji Arai.
It was a good surprise! Thanks, we will enjoy it!
I will return to Pasadena next week. See you then.
  665   Mon Jul 14 00:36:19 2008 JohnSummaryPSLSlow sweep of laser temp - PMC PZT response
John, Rana

Follow up to # 663

Top trace: C1: PSL-PMC_PZT
Middle: C1: PSL-FSS_SLOWDC
Bottom: C1: PSL-PMC_PMCTRANSPD

The only calibration I could find for the PMC PZT (LLO e-log Sep 3 2003 - 23 MHz/V) predicts 31V for an FSR. I did a rough calibration and got our FSR to be around 210 V. I assumed 713 MHz for an FSR and applied this calibration (~3.4 MHz/ V) to the PZT data.

In terms of volts per metre our PZT gives 2.54 nm/ V whereas the LLO PZT is 17.16 nm/ V.
Attachment 1: SlowTsweep.png
SlowTsweep.png
ELOG V3.1.3-